# Stable Distributed Online Feedback Optimization for Distribution System Voltage Regulation

### v1

@@ Abstract
We investigate the distributed voltage regulation problem in distribution systems employing online feedback optimization and short-range communication between physical neighbours. We show that a two-metric approach can be unstable. As a remedy, we propose a nested feedback optimization strategy. Simulation results reveal that while the two-metric approach fails to regulate voltages, the proposed approach achieves even less voltage limit violations than its centralized counterpart.
@@ I Introduction
With the proliferation of distributed energy resources (DERs) in distribution systems, voltage regulation has become a challenging task for distribution system operators (DSOs). While local control cannot guarantee the desired voltage regulation [1], model-based feedforward optimization requires accurate grid models, real-time load measurements, and lacks robustness to system uncertainties and disturbances [2].
Recently, online feedback optimization (OFO) has emerged as a promising approach for real-time distribution system control. OFO utilizes measurements as feedback and employs optimization algorithms as feedback controllers to steer the system towards an optimal operating point [2]. While many OFO controllers rely on centralized communication [3, 4], distributed OFO, which relies on communication between physical neighbours, appears to be a compelling approach due to its robustness to single-point failures [5, 1, 6, 7, 8].
While [5, 1, 6] demand particular objective function designs, [7] relies on disseminating global information through neighboring nodes. However, the systematic delays inherent in this information propagation may raise stability issues, potentially undermining its efficacy in practical distribution systems. In this letter, we propose a distributed OFO design that is closely related to that in [8], where the distributed communication is achieved by scaling the gradient with a sparse positive definite matrix $\mathbf{X}^{-1}$. A two-metric approach [9] was further used in [8], where an Euclidean projection step was used to ensure satisfaction of the DER capacity constraint. The approach is called two-metric because the matrix $\mathbf{X}^{-1}$ is used for the gradient scaling while the identity matrix $\mathbf{I}$ is used for the projection.
@@
However, utilizing such a two-metric approach might lead to divergence, as shown in [9]. Instead of the Euclidean distance, the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ should be used in the projection step to ensure convergence [9]. Solving this non-Euclidean projection problem would however require global information, contradicting the distributed OFO design. To tackle this challenge, we propose an iterative algorithm that requires only local information and operates in an online feedback-based manner. This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations yielding tentative but not necessarily feasible DER setpoints, and the inner loop solves the non-Euclidean projection to map the tentative setpoints to actual feasible setpoints.
@@ II-A Problem formulation
Consider the distribution system voltage regulation using reactive power as in Problem (1). DERs are located in the set $\mathcal{N}$, and $q_{i}$ represents their reactive power injections. The objective function (1a) minimizes the reactive power cost, where $c_{i}$ is the cost parameter. Constraint (1b) enforces the lower and upper voltage limits $\underline{v}$ and $\overline{v}$, where $\mu_{i}$ and $\lambda_{i}$ are the respective dual variables. Finally, (1c) ensures that the reactive power injection $q_{i}$ respects the lower and upper limits $\underline{q}_{i}$ and $\overline{q}_{i}$. Note that the voltage $v_{i}$ is dependent on $q_{i}$ and other disturbances in the system such as active power injections of the DERs and active and reactive power consumption of non-controllable load units.
@@ II-B Centralized primal-dual gradient projection algorithm
Solving Problem (1) offline in a feedforward way demands an accurate grid model and disturbance data to evaluate $v_{i}$, which might not be available. The feedforward approach also lacks robustness [2]. To tackle these challenges, OFO has emerged as a promising approach for real-time distribution system management. This section introduces an instance of OFO based on the primal-dual gradient projection (PDGP) algorithm [3], which requires a centralized gather-and-broadcast communication architecture. The PDGP algorithm performs projected gradient ascent and descent steps iteratively for the dual and primal variables, respectively. At each time step (iteration) $k$, it repeats the following three steps:
@@
    For each bus $i\in\mathcal{N}$, collect its voltage measurement $\tilde{v}_{i}^{k}$ and update $\lambda_{i}$ and $\mu_{i}$, where the projection operator is defined as $[u]^{+}:=\max(u,0)$ and $\alpha$ with various superscripts represents step sizes:
    $\lambda_{i}^{k+1}\leftarrow\big{[}\lambda_{i}^{k}+\alpha^{\lambda}(\tilde{v}_{i}^{k}-\overline{v})\big{]}^{+}$ (2a)
    $\mu_{i}^{k+1}\leftarrow\big{[}\mu_{i}^{k}+\alpha^{\mu}(\underline{v}-\tilde{v}_{i}^{k})\big{]}^{+}$ (2b)
@@
    For each bus $i\in\mathcal{N}$, update its reactive power setpoints $q_{i}$, where $[u]_{\underline{q}_{i}}^{\overline{q}_{i}}$ represents the Euclidean projection into $[\underline{q}_{i},\overline{q}_{i}]$, and $\mathbf{X}$ is a $|\mathcal{N}|$-dimensional dense matrix mapping reactive power injections to voltages, i.e., the sensitivity matrix, as constructed in [8]. Because of the non-sparsity of $\mathbf{X}$, Step 2 is required to provide information needed in Step 3.
    $q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[c_{i}q_{i}^{k}+[\mathbf{X}(\boldsymbol{\lambda}^{k+1}-\boldsymbol{\mu}^{k+1})]_{i}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (2c)
@@ II-C Two-metric distributed approach
To avoid a centralized communication architecture and promote the distributed communication architecture where communication is only between physical neighbours in the system, [8] proposed to scale the gradient in (2c) by a sparse positive definite matrix $\mathbf{X}^{-1}$, which has non-zero entries only in the diagonal and the $i$-th row $j$-th column if nodes $i$ and $j$ are adjacent. This yields (3) and replaces (2c) in the OFO, where $\mathbf{q}^{k}$ is the stack of $q^{k}_{i}$, and $\mathbf{C}$ is a $|\mathcal{N}|$-dimensional diagonal matrix with elements $c_{i}$ defined in (1a).
$q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (3)
Computing (3) for node $i$ now requires only information from its physcial neighbours. We stack $q_{i}^{k}-\alpha[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}$ for all nodes (tentative but not necessarily feasible setpoints to be projected) as the vector $\mathbf{\dot{q}}^{k+1}$ and define $\mathcal{X}^{k}$ as $[\underline{q}_{1}^{k},\overline{q}_{1}^{k}]\times[\underline{q}_{2}^{k},\overline{q}_{2}^{k}]\times\cdots\times[\underline{q}_{|\mathcal{N}|}^{k},\overline{q}_{|\mathcal{N}|}^{k}]$ for further developments, where $\times$ is the Cartesian product of sets.
@@ II-D Proposed distributed approach
The two-metric approach in Section II-C does not necessarily lead to algorithm convergence [9]. To ensure convergence, the projection in (3) should use the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ due to the use of the scaling matrix $\mathbf{X}^{-1}$ [9], that is:
@@
This however contradicts the distributed communication architecture of the overall OFO algorithm since directly solving (4) again requires global information. To tackle this challenge, this letter proposes an iterative algorithm that requires only local information and is implemented in an online feedback-based manner. This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations giving tentative reactive power setpoints, and the inner loop solves the non-Euclidean projection problem (4). Our key observation is that given a feasible $\mathbf{u}^{\tau}$, the gradient of the objective function in (4) is $\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})$, which can be approximated by the voltage difference between $\mathbf{v}(\mathbf{u}^{\tau})$ and $\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ assuming system disturbances do not change significantly when all nodes are locally implementing this inner iterative algorithm. The algorithm is based on the non-scaled Euclidean gradient projection and proceeds as:
@@
Note that these two steps can be implemented for each node using only local information. In the first step, $\epsilon$ represents a small exploration parameter, which is similar to that in model-free feedback optimization [10]. Since $\mathbf{\dot{q}}^{k+1}$ can be pretty far away from the feasible set, it might be impossible to implement it. Therefore, we implement $\mathbf{q}^{k}_{\epsilon}$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ based on it. This could still result in a minor temporary capacity limit violation, which should be manageable by the DER. To ensure strict capacity limit satisfaction, the original feasible set can be slightly deflated [10]. The initial feasible $\mathbf{u}^{\tau}$ in the second step can be simply chosen as $\mathbf{q}^{k}$.
@@ III Simulation results
In our simulation experiment, we compare the centralized OFO based on the gather-and-broadcast communication, the distributed two-metric approach in [8], and our proposed nested distributed approach in both static and dynamic cases. We study a 100% penetration scenario of photovoltaics (PVs) in a 96-node low-voltage distribution system, which experiences overvoltage issues. Details of the distribution system and the generation and load data with a 6-second resolution can be found in [11] and references therein. The PV inverters are assumed to be oversized by 20% to provide sufficient reactive power. The reactive power cost matrix $\mathbf{C}$ is chosen as the identity matrix. The lower and upper voltage limits are 0.95 and 1.05 pu, respectively. The unit for power is $kW/kVar$. The step sizes are chosen with a trial-and-error strategy [8] as $\alpha^{\lambda}=\alpha^{\mu}=10^{6}$, $\alpha=10^{-1}$ (for centralized), $\alpha=10^{-3}$ (for distributed), and $\alpha^{u}=10^{2}$. The exploration parameter $\epsilon$ is chosen as $10^{-5}$, resulting in a maximum DER reactive power limit violation of 2%. For each generation and load data point, an OFO iteration is run which yields tentative reactive power setpoints, followed by one exploration step and 5 inner feedback iterations of (5) ($T=5$).
@@
Figure 1: Voltage profiles over 200 iterations for two remote buses in the static case using the proposed and the two-metric approaches. We count both the outer and inner iterations for our proposed approach.
@@
Figure 2: Reactive power profiles after 200 iterations for the centralized, the proposed, and the two-metric approaches.
@@
Figure 3: Voltage profiles for the most remote bus in the dynamic simulation. The average voltage violation (AVV) values are $1.9\times 10^{-4}$, $1.3\times 10^{-4}$, and $4.6\times 10^{-3}$ pu for the centralized, the proposed, and the two-metric approaches, respectively, which is defined as $\operatorname{AVV}=\frac{1}{K}\sum_{k=1}^{K}[v^{k}-\overline{v}]^{+}$, where $k$ is the iteration index and $K$ is the number of iterations.
@@
In the static case, the generation and load data are kept unchanged to study the convergence of various approaches. Figure 1 shows that while the two-metric approach does not converge and fails to regulate voltages, our proposed approach converges quickly after approximately 154 iterations, which correspond to around 22 outer and 132 inner iterations. We also compare the final reached reactive power setpoints after 200 iterations. Figure 2 shows that the centralized and our proposed approaches converge to the same solution. This is reasonable since the gradient scaling only alters the trajectory, but does not change the solution to Problem (1).
@@
In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online without reaching convergence. Figure 3 shows the significant voltage limit violation without control and demonstrates that our proposed nested distributed approach successfully enforces the voltage limit even for the most remote, and therefore, most sensitive bus, achieving an average voltage violation (AVV) of $1.3\times 10^{-4}$ pu, which is better even than that of the centralized approach ($1.9\times 10^{-4}$ pu). We conjecture that this is because of the smoothing brought by the inner feedback loop. The two-metric approach, however, fails to regulate voltages, resulting in 34 times higher AVV.

### v2

# Distributed Online Feedback Optimization for Real-time Distribution System Voltage Regulation

@@ Abstract
We investigate the real-time distribution system voltage regulation problem employing online feedback optimization (OFO) and short-range communication between physical neighbours. OFO does not need an accurate grid topology nor estimated consumption of non-controllable loads, affords fast calculations, and demonstrates robustness to uncertainties due to its feedback-based nature. These characteristics render it particularly suitable for real-time distribution system applications. However, many OFO controllers require centralized communication, which lack robustness to single-point failures. This paper proposes a distributed OFO design based on a nested feedback optimization strategy and analyzes its convergence. Numerical results reveal that the proposed design achieves satisfactory voltage regulation and outperforms other distributed approaches.
@@ I Introduction
With the proliferation of distributed energy resources (DERs) in distribution systems, real-time voltage regulation has become a challenging task for distribution system operators (DSOs). While local control cannot guarantee the desired voltage regulation [1], model-based feedforward optimization requires an accurate grid model, load measurements, and lacks robustness to system uncertainties and disturbances [2].
To overcome these limitations, online feedback optimization (OFO) has been proposed as a promising solution approach. OFO utilizes measurements as feedback and employs optimization algorithms as feedback controllers to steer physical systems towards their optimal operating points [2]; see Fig. 1 for comparison between feedforward and feedback systems. The feedback-based implementation eliminates the need for an accurate grid model and real-time load data and brings robustness to OFO. Compared to deep reinforcement learning, OFO does not need a complicated training process and has guarantees to pursue grid-level objectives. Compared to using traditional voltage regulators such as on-load tap changers and capacitor banks, OFO provides fast voltage regulation and can complement these slow-reacting devices in real time [3].
While many OFO controllers rely on centralized communication [4, 5, 6, 7, 8], distributed OFO based on communication between physical neighbours appears a compelling approach due to its robustness to single-point failures[9, 1, 3, 10, 11, 12]. Among the existing distributed designs, [9, 1, 3] demand particular forms of objective functions, while [10] relies on disseminating global information through neighboring nodes. However, the systematic delays inherent in this information propagation may raise stability issues, potentially undermining its efficacy in practical distribution systems. In [11], the distributed design was achieved by only keeping components in the network sensitivity matrix related to local and neighbouring nodes. This however does not assure voltage regulation.
In this paper, we propose a distributed OFO design related to that in [12], where the distributed implementation was enabled by the sparsity of the inverse of the network sensitivity matrix $\mathbf{X}$, i.e. $\mathbf{X}^{-1}$ is sparse. Both approaches continue with scaling the gradient by $\mathbf{X}^{-1}$ followed by a projection step to ensure satisfaction of the DER capacity constraint. It is conceptually attractive to use the Euclidean distance in the projection, which can be implemented using only local information. However, as shown in [13], the distance measured by the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ should be used instead in the projection step to ensure descent iterations and thus algorithm convergence. With the Euclidean distance, it would be a two-metric approach which can lead to algorithm divergence [13]. Solving this non-Euclidean projection problem would however require global information since $\mathbf{X}$ is dense, contradicting the distributed OFO design. As a remedy, [12] introduced additional dualization of the DER capacity limits on top of the Euclidean projection to ensure the DER constraint satisfaction. Shown in their numerical result and our simulation, this strategy takes a prohibitively large number of iterations to converge, which jeopardizes its performance in online implementation.
Figure 1: Block diagrams of feedforward and feedback systems for voltage regulation using reactive power, where $q$ represents system input (reactive power), $v$ represents system output (voltages), $\omega$ ($\widehat{\omega}$) represents (estimated) system disturbance (non-controllable loads), and $h$ maps system input and disturbance to system output.
@@
We tackle this challenge by proposing an iterative algorithm to solve this non-Euclidean projection problem that requires only local information and operates in an online feedback-based manner. This leads to our overall nested online feedback optimization approach, where the outer loop comprises the OFO iterations requiring communication between physical neighbours and yielding tentative but not necessarily feasible DER setpoints, and the inner loop solves the non-Euclidean projection problem to map the tentative setpoints to actual feasible setpoints using local information.
The remainder of this paper is structured as follows: Section II presents the problem formulation and the proposed nested approach and analyzes its convergence. Section III presents simulation results, while Section IV draws conclusions and discusses future work.
@@ II-A Problem formulation
Consider the distribution system voltage regulation problem using reactive power as formulated in (1). DERs are located in the set $\mathcal{N}$, and $q_{i}$ represents their respective reactive power injection. The objective function (1a) minimizes the reactive power use cost, where $c_{i}>0$ is the cost parameter. Constraint (1b) enforces the lower and upper voltage limits $\underline{v}$ and $\overline{v}$, where $\mu_{i}$ and $\lambda_{i}$ are the respective dual variables. Finally, (1c) ensures that the reactive power injection $q_{i}$ respects the lower and upper limits $\underline{q}_{i}$ and $\overline{q}_{i}$. Note that $v_{i}$ is dependent on $q_{i}$ and other disturbances in the system such as active and reactive power consumption of non-controllable load units.
@@ II-B Centralized primal-dual gradient projection algorithm
Solving (1) offline in a feedforward way demands an accurate grid model and disturbance data to evaluate $v_{i}$, which might not be available. The feedforward approach also lacks robustness [2]. Thus, this section introduces an instance of OFO based on the primal-dual gradient projection (PDGP) algorithm [4, 5] to track the time-varying optimizer of (1), which requires a centralized gather-and-broadcast communication architecture. The PDGP algorithm performs projected gradient ascent and descent steps iteratively for the dual and primal variables, respectively. At each time step (iteration) $k$, it repeats the following three steps:
@@
    For each bus $i\in\mathcal{N}$, collect its voltage measurement $\tilde{v}_{i}^{k}$ and update $\lambda_{i}$ and $\mu_{i}$, where the projection operator is defined as $[u]^{+}:=\max(u,0)$, $\alpha$ with various superscripts represents step sizes, while $r^{d}$ and $r^{p}$ in (2c) are the dual and primal regularization factors [4, 5], respectively.
    $\lambda_{i}^{k+1}\leftarrow\big{[}\lambda_{i}^{k}+\alpha^{d}(\tilde{v}_{i}^{k}-\overline{v}-r^{d}\lambda_{i}^{k})\big{]}^{+}$ (2a)
    $\mu_{i}^{k+1}\leftarrow\big{[}\mu_{i}^{k}+\alpha^{d}(\underline{v}-\tilde{v}_{i}^{k}-r^{d}\mu_{i}^{k})\big{]}^{+}$ (2b)
@@
    For each bus $i\in\mathcal{N}$, update its reactive power setpoints $q_{i}$, where $[u]_{\underline{q}_{i}}^{\overline{q}_{i}}$ represents the Euclidean projection into $[\underline{q}_{i},\overline{q}_{i}]$, $\mathbf{X}$ is a $|\mathcal{N}|$-dimensional dense matrix mapping reactive power injections to voltages, i.e. the sensitivity matrix, as constructed in [12], and $\mathbf{q}^{k}$ is the stack of $q^{k}_{i}$. Since $\mathbf{X}$ is dense, Step 2 is needed to provide information for Step 3.
    $q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[c_{i}q_{i}^{k}+[\mathbf{X}(\boldsymbol{\lambda}^{k+1}-\boldsymbol{\mu}^{k+1}+r^{p}\mathbf{q}^{k})]_{i}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (2c)
@@ II-C Two-metric distributed approach
To avoid the centralized communication requirement and enable the distributed communication architecture where communication is only between physical neighbours, [12] proposed to scale the gradient in (2c) by a sparse positive definite matrix $\mathbf{X}^{-1}$, which has non-zero entries only in the diagonal and the $i$-th row $j$-th column if buses $i$ and $j$ are adjacent. Continuing with the Euclidean projection yields (3) which replaces (2c) in the OFO where $\mathbf{C}$ is a $|\mathcal{N}|$-dimensional diagonal matrix with $c_{i}$ defined in (1a).
$q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (3)
Computing (3) for node $i$ now requires only information exchange with its physical neighbours. We stack $q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}]$ for all nodes (tentative but not necessarily feasible setpoints to be projected) as the vector $\mathbf{\dot{q}}^{k+1}$ and define $\mathcal{X}^{k}$ as $[\underline{q}_{1}^{k},\overline{q}_{1}^{k}]\times[\underline{q}_{2}^{k},\overline{q}_{2}^{k}]\times\cdots\times[\underline{q}_{|\mathcal{N}|}^{k},\overline{q}_{|\mathcal{N}|}^{k}]$, where $\times$ is the Cartesian product of sets.
@@ II-D Proposed distributed approach
The two-metric approach in Section II-C does not necessarily lead to algorithm convergence since it is not in general a descent iteration [13]. To ensure convergence, the projection in (3) should use the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ due to the scaling matrix $\mathbf{X}^{-1}$ [13], that is:
@@
This however contradicts the distributed communication architecture of the overall OFO algorithm since, unlike the Euclidean projection (equivalent to the case where $\mathbf{X}$ is the identity matrix), solving (4) again requires global information. As a solution, [12] proposed dualization of the DER capacity limits, combined with the Euclidean projection, to ensure the DER constraint satisfaction. However, as demonstrated both in their numerical result and our simulation, this approach requires an excessively large number of iterations to converge, significantly hindering its performance in online implementation. To tackle this challenge, this paper proposes an iterative algorithm to solve (4) that requires only local information and is implemented in an online feedback-based manner. This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations giving tentative reactive power setpoints, and the inner loop solves the non-Euclidean projection problem (4) to generate feasible setpoints. Our key observation is that given a feasible $\mathbf{u}^{\tau}\in\mathcal{X}^{k}$, the gradient of the objective function in (4) at $\mathbf{u}^{\tau}$ is $\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})$, which can be approximated by the voltage difference $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ assuming system disturbances do not change significantly when all nodes are locally implementing this inner iterative algorithm. The proposed algorithm is based on the non-scaled Euclidean gradient projection and proceeds as:
@@
In the first step, $\epsilon$ represents a small exploration parameter, which is similar to that in model-free feedback optimization [14]. Since $\mathbf{\dot{q}}^{k+1}$ can be pretty far away from the feasible set, it might be impossible to implement it. Therefore, we implement $\mathbf{q}^{k}_{\epsilon}$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ based on it. This could still result in minor temporary DER capacity limit violations, which should be manageable by the DERs. To ensure strict capacity limit satisfaction, the original feasible set can be slightly deflated [14]. The initial feasible $\mathbf{u}^{\tau}$ in the second step can be chosen as $\mathbf{q}^{k}$. Figure 2 shows the implemented reactive power setpoints and communication requirements in an outer iteration.
Figure 2: Implemented reactive power setpoints and communication requirements in an outer iteration for the proposed strategy.
II-E Convergence analysis
The convergence analysis of the nested algorithm is based on the following two assumptions. First, we assume a linear relation holds between voltages and reactive power injections. That is, $\mathbf{v}=\mathbf{v}^{0}+\mathbf{X}\mathbf{q}$, where the constant term $\mathbf{v}^{0}$ captures the impact of other disturbances. Based on the first assumption, $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ is the exact gradient of the objective function in (4) at $\mathbf{u}^{\tau}$. The inner loop then represents a standard Euclidean gradient projection algorithm to solve the quadratic program (4). Consequently, $\mathbf{u}^{\tau}$} is a converging sequence by invoking Proposition 2.3.2 in [13] and $\alpha^{u}$ chosen with the rule therein. Second, we assume the inner iterations indeed converge after $T$ iterations. The overall nested algorithm thus represents a scaled PDGP algorithm to solve (1). By a transformation of variables defined by: $\mathbf{q}^{\prime}=\mathbf{X}^{\frac{1}{2}}\mathbf{q}$ and $\mathbf{v}^{\prime}=\gamma\mathbf{v}$ where $\gamma=\frac{\alpha^{d}}{\alpha}$ is the ratio between the dual and primal step sizes, the problem (1) can be cast as a quadratic program in the space of $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$. Likewise, the overall nested algorithm can be cast as an equivalent unscaled PDGP algorithm in $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$ [13]. Invoking Theorem 3 in [5] and with $\alpha$ chosen sufficiently small within the upper bound defined therein, the overall algorithm then converges to the optimizer of the regularized saddle-point problem of (1) at a Q-linear rate.
Remark: First, while the linear voltage-reactive power relation is assumed to synthesize the controller and analyze its convergence, AC power flow is run in the simulation to calculate the real voltages. Such an assumption was also made in the proofs of [1, 11]. Second, a large $T$ is not necessary in the simulation. We use $T=4$ so the inner loop might not converge. However, the algorithm performs well in the simulation. Notably, a similar assumption was made in [1].
@@ III Simulation results
In the numerical simulation, we study a 100% penetration scenario of photovoltaics (PVs) in a 96-node system, which experiences overvoltage issues. Details of the system and the generation and load data with a 6-second resolution can be found in [7] and references therein. The PV inverters are assumed to be oversized by 20% to provide sufficient reactive power. The reactive power cost matrix $\mathbf{C}$ is chosen as the identity matrix. The lower and upper voltage limits are 0.95 and 1.05 pu, respectively. The unit for power is $kW/kVar$. The step sizes are chosen with a trial-and-error strategy [12] as $\alpha^{d}=10^{6}$, $\alpha=2\times 10^{-4}$, and $\alpha^{u}=10^{2}$. The exploration parameter $\epsilon$ is chosen as $10^{-5}$, resulting in a maximum DER reactive power limit violation of 1%. The primal and dual regularization is used to establish the theoretical convergence result. The impact of some small regularization parameters on the numerical results is negligible. In the simulation, we set them as $10^{-4}$. For each generation and load data point, an OFO iteration is run which yields tentative reactive power setpoints, followed by one exploration step and 4 inner feedback iterations of (5), i.e. we assume that every setpoint in Fig. 2 is implemented for 1 second. Finally, the nested algorithm requires only basic arithmetic operations and is computationally very efficient. For a series implementation of a 4-hour simulation, i.e. in total 14400 iterations, it takes 7.6 seconds, averaging 0.5 millisecond per iteration.
@@
Figure 3: Voltage profiles over 1000 iterations for several buses in the static case using the centralized and the proposed approaches. We count both the outer and inner iterations for our proposed approach.
@@
Figure 4: Reactive power profiles after 1000 iterations for the centralized and the proposed approaches.
@@
Figure 5: Voltage profiles for the most remote bus in the dynamic simulation. We define the average voltage violation (AVV) as $\operatorname{AVV}=\frac{1}{K}\sum_{k=1}^{K}[v^{k}-\overline{v}]^{+}$, where $k$ is the iteration index and $K$ is the number of iterations. The AVV values for the various approaches are: $1.5\times 10^{-2}$ (no control), $8.3\times 10^{-5}$ (centralized), $2.5\times 10^{-3}$ (two-metric), $2.1\times 10^{-4}$ (proposed), $3.3\times 10^{-3}$ (Tang 2024 [11]), $8.6\times 10^{-4}$ (Qu 2020 [12]).
@@
In the static case, the generation and load data are kept unchanged to study the convergence of the nested approach. Figure 3 shows that the proposed approach suppresses the initial voltage limit violations quickly and converges after around 300 total iterations, i.e. 50 outer OFO iterations. The approach also utilizes the available grid capacity as efficiently as the centralized one. Figure 4 compares the converged reactive power setpoints, which shows that the centralized and the proposed approaches converge to the same solution. This is reasonable since the gradient scaling only alters the converging trajectory, but does not change the solution to (1).
@@
In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online to track the time-varying optimizer of (1) without waiting for them to converge. Figure 5 shows significant voltage limit violations without control and demonstrates that our proposed distributed nested feedback optimization approach successfully enforces the voltage limit even for the most remote, and therefore, the most sensitive bus. An average voltage violation (AVV), defined in Fig. 5, of $2.1\times 10^{-4}$ pu is achieved, which is only higher than that of the centralized approach. The latter however requires a centralized communication architecture and is thus susceptible to single-point failures. The proposed nested feedback approach significantly outperforms other distributed ones, which is visible in Fig. 5 and quantified using the AVV.